OpenStudy (anonymous):

what is the least common denominator of 2,3,and 15

OpenStudy (anonymous):

since 15 is an odd number and it must share a common denominator with 2, we know it will be at least 30 (2*15)

OpenStudy (anonymous):

we know 15=3*5, so 30 will work for 3 as well

OpenStudy (anonymous):

so 30 is our LCD 2*15=30 3*10=30 15*2=30